# Optimistic Query Routing in Clustering-based Approximate Maximum Inner Product Search

### v1

@@ Appendix E Effect of parameters
Figure 9: Top-$100$ recall versus the number of probed data points. Datasets are partitioned with Spherical KMeans. We compare a range of values for the rank of the covariance sketch ($t$) with SubPartition with the same capacity as Optimist with the smallest and largest ranks. $\textsc{Optimist}(0,\cdot)$ denotes a sketch that only retains the diagonal of the covariance matrix; and $\textsc{SubPartition}(t)$ means every shard is partitioned into $t+2$ sub-partitions.
@@
Figure 10: Top-$100$ recall versus the number of probed data points. Datasets are partitioned with Spherical KMeans. We compare a range of values for the optimism parameter ($\delta$) with the strongest baseline (SubPartition) from Appendix D.1. As $\delta\rightarrow 1$, Optimist becomes more optimistic.
Appendix F Prediction error
@@
Figure 11: Mean prediction error $\mathcal{E}_{\ell}(\tau,\cdot)$, defined in Equation (8), versus $\ell$ (expressed as percent of total number of shards), for various routers and datasets.

### v3

@@ 3.3. Effect of parameters on Optimist
Figure 6. Top-$100$ recall vs. volume of probed data as we change the rank of the covariance sketch ($t$). $\textsc{Optimist}(0,\cdot)$ denotes a sketch that only retains the diagonals. Partitioning is with Spherical KMeans.
Recall that Optimist takes two parameters: $t$, the rank of the covariance sketch, and $\delta$, the degree of optimism. We examine the effect of these parameters on the performance of Optimist.
In the same figure, we have also included two configurations of SubPartition: one with $2$ sub-partitions, $\textsc{SubPartition}(0)$, and another with $t+2$ sub-partitions, $\textsc{SubPartition}(t)$, for the largest $t$. These help put the performance of Optimist with various ranks in perspective. In particular, we give the SubPartition baseline the same amount of information and contrast its recall with Optimist.
@@
Figure 7. Top-$100$ recall vs. volume of probed data, comparing a range of values of $\delta$. As $\delta\rightarrow 1$, Optimist becomes more optimistic. Partitioning is with Spherical KMeans.
3.4. Maximum inner product prediction
@@
Figure 8. Mean prediction error $\mathcal{E}_{\ell}(\tau,\cdot)$, defined in Equation (24), versus $\ell$ (expressed as percent of total number of shards), for various routers and datasets.
We claimed that Optimist is statistically principled. We expect then, that it should give more accurate estimates of the maximum inner product for any given query-partition pair. We examine that claim in this section and quantify the prediction error for all routers.
Fix a dataset, whose $C$ partitions are denoted by $\mathcal{P}_{i}$ for $i\in[C]$, along with a router $\tau$ and query $q$. We write $\tau_{i}$ to denote the score computed by $\tau$ for partition $\mathcal{P}_{i}$ and $q$; for example, NormalizedMean computes $\langle q,\mu_{i}/\lVert\mu_{i}\rVert_{2}\rangle$ and Optimist $\theta_{i}$ per Equation (6).
(24) $\mathcal{E}_{\ell}(\tau,q)=\frac{1}{\ell}\sum_{i=1}^{\ell}\left|\frac{\tau_{\pi_{i}}}{\max_{u\in\mathcal{P}_{\pi_{i}}}\langle q,u\rangle}-1\right|.$
We measure Equation (24) on all datasets partitioned by Spherical KMeans, and all routers considered in this work. The results are shown in Figure 8, where for each choice of $\ell$, we plot $\mathbb{E}_{q}[\mathcal{E}_{\ell}(\tau,q)]$ for test queries.
[…]
4. Concluding remarks
We studied clustering-based maximum inner product search, where data points are clustered into non-overlapping shards to form an index. At query time, two independent components work together to process a query: a router computes a score for each shard with respect to the query and identifies the subset of top shards to probe; and a subsequent scoring stage that computes (approximate) scores for points in the chosen shards. In this general setup, we considered a storage-backed system where shards do not rest in the main memory, but are stored on external storage and must be fetched into the main memory when a router identifies them. This is an important paradigm for vector search at scale.
Within this framework, we focused on routing. We motivated our work by an unusual but unsurprising observation: NormalizedMean computes a score for each shard that over-estimates the maximum inner product between an arbitrary query and the points in that shard. Interestingly, the extent of over-estimation correlates with the variance of the shard: The more spread-out the points are, the more optimistic the NormalizedMean router becomes.
This behavior is unpredictable, however. It is also hard to explain why it leads to more accurate routing when data points have non-constant norms. But we found the insight that variance should play a role in routing to be worth investigating. We formalized that insight and presented a principled optimistic algorithm whose score for a shard is a more accurate estimate of the maximum inner product—as confirmed by evidence in Figure 8.
As Figure 5 shows, our algorithm, Optimist, is more suitable for settings where shards rest on some external, high-latency storage. That is because, in such settings, spending more computation on routing is acceptable if we end up transferring a smaller amount of data from storage to main memory. It is particularly attractive when the storage medium is a cheap, high-latency blob storage.
Optimist has other implications beyond latency. Needing to transfer a smaller volume of data to memory means that each query requires a smaller chunk of the in-memory cache. As a result, cache can be shared among more queries, leading to a higher throughput in systems where memory availability is the main bottleneck.
Our research takes a first step in exploring unsupervised routing for clustering-based MIPS, and identifies a trade-off space that has not been explored previously. We have more to do, however. We leave to future work an exploration of a more compact sketch of the covariance matrix; and, an efficient realization of our general solution outlined in Section 2.3.
Appendix A Additional experimental results
(a) Nq-Ada2
(b) Music
(c) GloVe
Figure 9. Top-$k$ recall vs. volume of probed data. Partitioning is with Standard KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
(d) Text2Image
(e) MsMarco-MiniLM
(f) DeepImage
Figure 9. Top-$k$ recall vs. volume of probed data. Partitioning is with Standard KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
(a) Music
(b) GloVe
(c) Text2Image
Figure 10. Top-$k$ recall vs. volume of probed data. Partitioning is with Gaussian Mixture Model. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$. We note that, due to the dimensionality of Nq-Ada2, we were unable to complete GMM clustering on this particular dataset.
(d) MsMarco-MiniLM
(e) DeepImage
Figure 10. Top-$k$ recall vs. volume of probed data. Partitioning is with Gaussian Mixture Model. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$. We note that, due to the dimensionality of Nq-Ada2, we were unable to complete GMM clustering on this particular dataset.